# Generalization Measures for Zero-Shot Cross-Lingual Transfer

### v1

@@ Abstract
A model’s capacity to generalize its knowledge to interpret unseen inputs with different characteristics is crucial to build robust and reliable machine learning systems. Language model evaluation tasks lack information metrics about model generalization and their applicability in a new setting is measured using task and language-specific downstream performance, which is often lacking in many languages and tasks. In this paper, we explore a set of efficient and reliable measures that could aid in computing more information related to the generalization capability of language models in cross-lingual zero-shot settings. In addition to traditional measures such as variance in parameters after training and distance from initialization, we also measure the effectiveness of sharpness in loss landscape in capturing the success in cross-lingual transfer and propose a novel and stable algorithm to reliably compute the sharpness of a model optimum that correlates to generalization.  [^1] Code: https://anonymous.4open.science/r/strikegen-7288
@@ 1 Introduction
Generalization enables models to use prior knowledge to reasonably respond to previously unseen stimuli. Although traditional machine learning evaluation is performed based on a preselected set of prediction or generation tasks, accuracy on many public benchmarks may often not be sufficient to extensively assess the ability to perform well in new settings. Therefore, a majority of researchers have found it worthwhile to investigate measures that could evaluate the generalization capability of models with properties, such as VC dimension Vapnik & Chervonenkis (1971), cross-entropy Shannon (1948), complexity Mohri et al. (2012) or variation in parameters during training Nagarajan & Kolter (2019). Among these, recent findings support the smoothness in the loss curvature to correlate best with generalization capability Chaudhari et al. (2019); Petzka et al. (2021); Kaddour et al. (2022), motivating the development of learning methods that induce smoothness in the learning trajectory such that the model becomes more robust; either through data perturbation Jiang et al. (2020a); Aghajanyan et al. (2021); Liang et al. (2021); Hua et al. (2021); Park et al. (2022); Zheng et al. (2021); Wang et al. (2021); Huang et al. (2021) or by integrating the measure directly to the optimization objective Izmailov et al. (2018); Jastrzebski et al. (2021); Cha et al. (2021); Foret et al. (2021); Hu et al. (2022); Zaken et al. (2022); Stickland & Murray (2021). However it might often not be straightforward to compute such measures in high-dimensional feature space in a stable fashion Nachum et al. (2024).
@@
  • We pick prominent measures that were previously shown to correlate well with generalization performance Jiang et al. (2020b), such as the Frobenius distance of the learned parameters after training Nagarajan & Kolter (2019), the margin between model predictions and true labels Wei et al. (2018) and sharpness in loss minima to test applicability to zero-shot cross-lingual generalization measurement Keskar et al. (2017); Foret et al. (2021).
@@ 2 Related work
Loss-landscape Minima One of the most promising indicators of generalization capability to date seems to be related to the form of the loss landscape, in particular, the sharpness in the loss curvature. A potential reason for this fallback is traced to stochastic gradient descent (SGD) Bottou (2012) methods which often fall into sharp minima of the loss surface Keskar et al. (2017); Chaudhari et al. (2019); Wang et al. (2021). Although clear conclusions on the relationship between sharpness and generalization performance, such as whether sharper Dinh et al. (2017) vs. flatter Li et al. (2018); Keskar et al. (2017) minima would generally yield better generalization, are still due. The main idea behind these methods is that their objective is to explicitly find flat minima, often using stochastic averaging methods Polyak & Juditsky (1992); Izmailov et al. (2018), mini-max or sharpness-aware minimization methods, which can be computed by direct formulation based on the Hessian matrix of the loss function Chaudhari et al. (2019); Petzka et al. (2021) or Monte-Carlo approximations of the minimizer’s neighborhood Foret et al. (2021); Cha et al. (2021).
Adversarial optimization Comparison of two approaches finds that for NLP tasks, mini-max methods are more competitive over averaging-based optimization Kaddour et al. (2022). Jastrzebski et al. (2021) hypothesize that regularizing the trace of the Fisher information matrix amplifies the implicit bias of SGD, which prevents memorization. The Fisher information Fisher (1925) measures local curvature, so a smaller trace implies a flatter minimum, which gives the model more freedom to reach an optimum. Instead of explicitly minimizing the values of parameters, Foret et al. (2021) propose minimizing both loss and sharpness while optimizing the parameters such that they lie in neighborhoods with low loss values. Perturbation is an auxiliary objective that encourages the model predictions to be similar in the vicinity of the observed training samples Englesson & Azizpour (2021), usually by penalizing the KL-divergence between the probability distribution of the perturbed and normal model. Perturbations can be adversarial inputs Jiang et al. (2020a) or inputs with Gaussian or uniform noise Aghajanyan et al. (2021). To improve cross-lingual generalization, translations of the input generated by machine translation systems were used as perturbed input Wang et al. (2021); Zheng et al. (2021). Other work also has found the benefit of enforcing consistency for perturbations within the model in addition to the input distribution Liang et al. (2021); Hua et al. (2021).
@@ 3.1 Sharpness-based Optimization
We chose the following objective functions as fine-tuning methods for a given pre-trained model as a means of comparison since their main purpose is to enhance the generalization and robustness of models. Following the work of Stickland & Murray (2021), as the two most prominent approaches to mini-max optimization, we include Sharpness-Aware Minimization (SAM) Foret et al. (2021) and regularization with Fisher Information Matrix (FIM) Jastrzebski et al. (2021) in our evaluation study on cross-lingual generalization.We also include Multi-view Subword Regularization (MVR) as a perturbation-based optimization method Wang et al. (2021) which induces stochasticity into the shared subword vocabulary across languages for easing cross-lingual transfer.
SAM Foret et al. (2021) works on the principle of a mini-max objective function: $\min_{w}\max_{\|\epsilon\|_{2}<\rho}L(w+\epsilon)$, which essentially means the optimizing function tries to minimize the maximum loss value in a given radius in loss landscape. Therefore, SAM states that it tries to seek ”parameters lying in uniformly low-loss neighborhoods”.
Fisher Penalty is defined as explicitly penalizing the trace of the Fisher information matrix (FIM). Jastrzebski et al. (2021), Stickland & Murray (2021) observed penalizing FIM during training correlates to better generalization performance. It can be written mathematically as $\frac{1}{n}\sum_{i}^{n}\nabla L(x_{i},y_{i})$ where $L(x_{i},y_{i})$ is the loss at the data point $(x_{i},y_{i})$.
@@ 4.1 Data, Model details, and Settings
For this work, we used the XNLI dataset Conneau et al. (2018) that includes data samples from the MultiNLI dataset Williams et al. (2018) and their translated versions in 14 different languages (Arabic ”ar”, Bulgarian ”bg”, German ”de”, Greek ”el”, Spanish ”es”, French ”fr”, Hindi ”hi”, Russian ”ru”, Swahili ”sw”, Thai ”th”, Turkish ”tr”, Urdu ”ur”, Vietnamese ”vi”, Chinese ”zh”). We only train the models on the English (”en”) subset of the dataset. We use the data of these 14 languages only for inference and evaluation of the models.
We fine-tune pretrained mBERT models for 15 epochs each with a batch size of 32, with a learning rate of $2\times 10^{-5}$, and select best checkpoint on validation. The objective function we use for the baseline model with the classification layer is the AdamW optimizer Loshchilov & Hutter (2019) with cross-entropy loss, the mBERT+FIM model has an additional loss as Fisher Penalty, the mBERT+SAM model uses the SAM optimizer and mBERT+MVR uses the MVR algorithm for fine-tuning. We use the hyperparameters and code presented in XTREME [^2] https://github.com/google-research/xtreme and MVR codebase [^3] https://github.com/cindyxinyiwang/multiview-subword-regularization. We run the models with 8 random seeds and present the average performance of these models (Figure 6). In Algorithm 1, the amount of Gaussian noise we add to model weights during calculating sharpness is controlled using a scale that we empirically find (among [0.001, 0.005, 0.01, 0.02]) for each model, with $n$ equal to 0.05.
@@
Additional experiments were run on PAWS-X dataset Yang et al. (2019) which has 7 languages: German ”de”, English ”en”, Spanish ”es”, French ”fr”, Japanese ”ja”, Korean ”ko”, Chinese ”zh”. We use similar experimentation of fine-tuning on english and doing a zero-shot transfer on 6 other languages as defined above for this dataset. We used Huggingface’s models: mBERT (bert-base-multilingual-cased), RoBERTa (roberta-base), and XLM (xlm-mlm-en-2048) using Adam optimizers.
@@ Results
We notice similar results by extending our similar experimentation to Paraphrase Identification, PAWS-X dataset Yang et al. (2019) with 3 different models: mBERT (bert-base-multilingual-cased), RoBERTa (roberta-base) Liu et al. (2019), and XLM (xlm-mlm-en-2048) CONNEAU & Lample (2019) and analyze the validity of the flatness hypothesis, i.e. a flat optimum neighborhood would lead to a generalized model. In Figure 1, we confirm the strong relationship between Margin (indicating generalization) and Sharpness (indicating flatness) even when compared across all models and metrics, suggesting flat neighborhoods of model optimum can help in achieving higher margin values which correlate to better generalization. More findings about visualizations are in Appendix A.1.
@@
5 Conclusion
6 Limitations
Appendix A Appendix
A.1 Visualization results
@@
Previous work Nagarajan & Kolter (2019); Jiang et al. (2020b) suggests that a lower Frobenius distance from initialization would lead to better generalization. As Figure 5 shows, we fail to observe a strong direct relationship between generalization and Frobenius distance from initialization. However, the model trained with Fisher Penalty as an additional objective function that has a high distance from initialization overall performed poorly than others. We also see that models trained with Fisher Penalty, SAM, and MVR optimizers tend to be more stable than the baseline model, with Fisher Penalty resulting in the most stable model when trained multiple times (with different seeds, see Figure 6), and SAM achieving generally the best average zero-shot task accuracy.
@@ A.2 Additional experiments
Model Correlation coefficient of $\alpha$ sharpness with accuracy

### v2

@@ Abstract
Building robust and reliable machine learning systems requires models with the capacity to generalize their knowledge to interpret unseen inputs with different characteristics. Traditional language model evaluation tasks lack informative metrics about model generalization, and their applicability in new settings is often measured using task and language-specific downstream performance, which is lacking in many languages and tasks. To address this gap, we explore a set of efficient and reliable measures that could aid in computing more information related to the generalization capability of language models, particularly in cross-lingual zero-shot settings. Our central hypothesis is that the sharpness of a model’s loss landscape, i.e., the representation of loss values over its weight space, can indicate its generalization potential, with a flatter landscape suggesting better generalization. We propose a novel and stable algorithm to reliably compute the sharpness of a model optimum, and demonstrate its correlation with successful cross-lingual transfer.  [^1] Code: https://anonymous.4open.science/r/strikegen-7288
@@ 1 Introduction
Generalization enables models to use prior knowledge to reasonably respond to previously unseen stimuli. Although traditional machine learning evaluation is performed based on a preselected set of prediction or generation tasks, accuracy on many public benchmarks may often not be sufficient to extensively assess the ability to perform well in new settings. Therefore, a majority of researchers have found it worthwhile to investigate measures that could evaluate the generalization capability of models with properties, such as VC dimension Vapnik and Chervonenkis (1971), cross-entropy Shannon (1948), complexity Mohri et al. (2012) or variation in parameters during training Nagarajan and Kolter (2019). Among these, recent findings support the smoothness in the loss curvature to correlate best with generalization capability Chaudhari et al. (2019); Petzka et al. (2021); Kaddour et al. (2022), motivating the development of learning methods that induce smoothness in the learning trajectory such that the model becomes more robust; either through data perturbation Jiang et al. (2020a); Aghajanyan et al. (2021); Liang et al. (2021); Hua et al. (2021); Park et al. (2022); Zheng et al. (2021); Wang et al. (2021); Huang et al. (2021) or by integrating the measure directly to the optimization objective Izmailov et al. (2018); Jastrzebski et al. (2021); Cha et al. (2021); Foret et al. (2021); Hu et al. (2022); Zaken et al. (2022); Stickland and Murray (2021). However it might often not be straightforward to compute such measures in high-dimensional feature space in a stable fashion Nachum et al. (2024).
@@
  • We pick prominent measures that were previously shown to correlate well with generalization performance Jiang et al. (2020b), such as the Frobenius distance of the learned parameters after training Nagarajan and Kolter (2019), the margin between model predictions and true labels Wei et al. (2018) and sharpness in loss minima to test applicability to zero-shot cross-lingual generalization measurement Keskar et al. (2017); Foret et al. (2021).
@@ 2 Related work
Loss-landscape Minima One of the most promising indicators of generalization capability to date seems to be related to the form of the loss landscape, in particular, the sharpness in the loss curvature. A potential reason for this fallback is traced to stochastic gradient descent (SGD) Bottou (2012) methods which often fall into sharp minima of the loss surface Keskar et al. (2017); Chaudhari et al. (2019); Wang et al. (2021). Although clear conclusions on the relationship between sharpness and generalization performance, such as whether sharper Dinh et al. (2017) vs. flatter Li et al. (2018); Keskar et al. (2017) minima would generally yield better generalization, are still due. The main idea behind these methods is that their objective is to explicitly find flat minima, often using stochastic averaging methods Polyak and Juditsky (1992); Izmailov et al. (2018), mini-max or sharpness-aware minimization methods, which can be computed by direct formulation based on the Hessian matrix of the loss function Chaudhari et al. (2019); Petzka et al. (2021) or Monte-Carlo approximations of the minimizer’s neighborhood Foret et al. (2021); Cha et al. (2021).
Adversarial optimization Comparison of two approaches finds that for NLP tasks, mini-max methods are more competitive over averaging-based optimization Kaddour et al. (2022). Jastrzebski et al. (2021) hypothesize that regularizing the trace of the Fisher information matrix amplifies the implicit bias of SGD, which prevents memorization. The Fisher information Fisher (1925) measures local curvature, so a smaller trace implies a flatter minimum, which gives the model more freedom to reach an optimum. Instead of explicitly minimizing the values of parameters, Foret et al. (2021) propose minimizing both loss and sharpness while optimizing the parameters such that they lie in neighborhoods with low loss values. Perturbation is an auxiliary objective that encourages the model predictions to be similar in the vicinity of the observed training samples Englesson and Azizpour (2021), usually by penalizing the KL-divergence between the probability distribution of the perturbed and normal model. Perturbations can be adversarial inputs Jiang et al. (2020a) or inputs with Gaussian or uniform noise Aghajanyan et al. (2021). To improve cross-lingual generalization, translations of the input generated by machine translation systems were used as perturbed input Wang et al. (2021); Zheng et al. (2021). Other work also has found the benefit of enforcing consistency for perturbations within the model in addition to the input distribution Liang et al. (2021); Hua et al. (2021).
@@ 3.1 Sharpness-based Optimization
We chose the following objective functions as fine-tuning methods for a given pre-trained model as a means of comparison since their main purpose is to enhance the generalization and robustness of models. Following the work of Stickland and Murray (2021), as the two most prominent approaches to mini-max optimization, we include Sharpness-Aware Minimization (SAM) Foret et al. (2021) and regularization with Fisher Information Matrix (FIM) Jastrzebski et al. (2021) in our evaluation study on cross-lingual generalization.We also include Multi-view Subword Regularization (MVR) as a perturbation-based optimization method Wang et al. (2021) which induces stochasticity into the shared subword vocabulary across languages for easing cross-lingual transfer.
SAM Foret et al. (2021) works on the principle of a mini-max objective function: $\min_{w}\max_{\|\epsilon\|_{2}<\rho}L(w+\epsilon)$, which essentially means the optimizing function tries to minimize the maximum loss value in a given radius in loss landscape. Therefore, SAM states that it tries to seek "parameters lying in uniformly low-loss neighborhoods".
Fisher Penalty is defined as explicitly penalizing the trace of the Fisher information matrix (FIM). Jastrzebski et al. (2021), Stickland and Murray (2021) observed penalizing FIM during training correlates to better generalization performance. It can be written mathematically as $\frac{1}{n}\sum_{i}^{n}\nabla L(x_{i},y_{i})$ where $L(x_{i},y_{i})$ is the loss at the data point $(x_{i},y_{i})$.
@@ 4.1 Data, Model details, and Settings
For this work, we used the XNLI dataset Conneau et al. (2018) that includes data samples from the MultiNLI dataset Williams et al. (2018) and their translated versions in 14 different languages (Arabic "ar", Bulgarian "bg", German "de", Greek "el", Spanish "es", French "fr", Hindi "hi", Russian "ru", Swahili "sw", Thai "th", Turkish "tr", Urdu "ur", Vietnamese "vi", Chinese "zh"). We only train the models on the English ("en") subset of the dataset. We use the data of these 14 languages only for inference and evaluation of the models.
We fine-tune pretrained mBERT models for 15 epochs each with a batch size of 32, with a learning rate of $2\times 10^{-5}$, and select best checkpoint on validation. The objective function we use for the baseline model with the classification layer is the AdamW optimizer Loshchilov and Hutter (2019) with cross-entropy loss, the mBERT+FIM model has an additional loss as Fisher Penalty, the mBERT+SAM model uses the SAM optimizer and mBERT+MVR uses the MVR algorithm for fine-tuning. We use the hyperparameters and code presented in XTREME [^2] https://github.com/google-research/xtreme and MVR codebase [^3] https://github.com/cindyxinyiwang/multiview-subword-regularization. We run the models with 8 random seeds and present the average performance of these models (Figure 6). In Algorithm 1, the amount of Gaussian noise we add to model weights during calculating sharpness is controlled using a scale that we empirically find (among [0.001, 0.005, 0.01, 0.02]) for each model, with $n$ equal to 0.05.
@@
Additional experiments were run on PAWS-X dataset Yang et al. (2019) which has 7 languages: German "de", English "en", Spanish "es", French "fr", Japanese "ja", Korean "ko", Chinese "zh". We use similar experimentation of fine-tuning on english and doing a zero-shot transfer on 6 other languages as defined above for this dataset. We used Huggingface’s models: mBERT (bert-base-multilingual-cased), RoBERTa (roberta-base), and XLM (xlm-mlm-en-2048) using Adam optimizers.
@@ Results
We notice similar results by extending our similar experimentation to Paraphrase Identification, PAWS-X dataset Yang et al. (2019) with 3 different models: mBERT (bert-base-multilingual-cased), RoBERTa (roberta-base) Liu et al. (2019), and XLM (xlm-mlm-en-2048) CONNEAU and Lample (2019) and analyze the validity of the flatness hypothesis, i.e. a flat optimum neighborhood would lead to a generalized model. In Figure 1, we confirm the strong relationship between Margin (indicating generalization) and Sharpness (indicating flatness) even when compared across all models and metrics, suggesting flat neighborhoods of model optimum can help in achieving higher margin values which correlate to better generalization. More findings about visualizations are in Appendix A.1.
@@
5 Conclusion
Limitations
Appendix A Appendix
A.1 Visualization results
Previous work Nagarajan and Kolter (2019); Jiang et al. (2020b) suggests that a lower Frobenius distance from initialization would lead to better generalization. As Figure 5 shows, we fail to observe a strong direct relationship between generalization and Frobenius distance from initialization. However, the model trained with Fisher Penalty as an additional objective function that has a high distance from initialization overall performed poorly than others. We also see that models trained with Fisher Penalty, SAM, and MVR optimizers tend to be more stable than the baseline model, with Fisher Penalty resulting in the most stable model when trained multiple times (with different seeds, see Figure 6), and SAM achieving generally the best average zero-shot task accuracy.
@@ A.2 Additional experiments
Model Correlation coefficient of
$\alpha$-sharpness with accuracy